a capacior of 2micro faraday is cinected to a tank circuit of an oscillator oscillating with a frequency 1kilohertz if current flowing in the circuit is 2milliamp the volyage across the capacitor will be

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
V_C=iX_C=2\times 10^{-3}\times \dfrac{1}{4\pi}\times 10^6V=159.15VHence voltage across the capacitor=
 
Hope it helps
In case of any difficulty, please feel free to ask.
 
Regards
Arun